OpenStudy (anonymous):

Ismael is comparing cell phone plans before upgrading his phone. Ameri-Mobile offers a low activation fee, but a high monthly payment. Cell-U-Later offers a lower monthly rate, but the activation fee is higher. Create a possible algebraic expression for both Ameri-Mobile and Cell-U-Later that shows the amount paid after an unknown amount of months have passed. Justify how you created those expressions and identify what each term and factor represents in terms of the cell phone plans.

OpenStudy (anonymous):

heyi have thisdou want the awnser??@WhiteFlower

OpenStudy (anonymous):

AM-20+100x am=ameri-mobile (af=activation fee) (MP=monthly payment) CUL+300-50x (CUL=Cell-U-Later)(xM=unknown months)i created the variables by using the initials of the words thats a equation that i used on the 01.04 quiz

OpenStudy (anonymous):

I just figured it out a few minutes ago :) Thanks though! Since you answered Ill give you the medal (:

OpenStudy (anonymous):

thankss message me if u need anymore help on the quiz @WhiteFlower